The drawing below shows a square with side a. A straight line intersects the square and encloses

an area A. The heights x and y on the left and right side (in a distance d from the square) of

the intersecting line can be varied. Assuming that x  y and x; y  a, nd an expression for

the enclosed area A(x; y) with respect to x and y.



1
Expert's answer
2021-12-14T05:47:58-0500


Area of the trapezoid "BCDE"


"A=\dfrac{BE+CD}{2}\cdot ED"

"\dfrac{BM}{FM}=\dfrac{CN}{FN}=\dfrac{GK}{FK}"

Substitute


"\dfrac{BE-x}{d}=\dfrac{CD-x}{a+d}=\dfrac{y-x}{a+2d}"

"d(BE-x)+a(BE-x)=d(CD-x)"

"d=\dfrac{a(BE-x)}{CD-BE}"

"a+d=\dfrac{a(CD-BE+BE-x)}{CD-BE}=\dfrac{a(CD-x)}{CD-BE}"

"a+2d=\dfrac{a(CD-BE+2BE-2x)}{CD-BE}"

"=\dfrac{a(CD+BE-2x)}{CD-BE}"

"\dfrac{BE-x}{d}=\dfrac{CD-BE}{a}"

"\dfrac{y-x}{a+2d}=\dfrac{(y-x)(CD-BE)}{a(CD+BE-2x)}"

Then


"\dfrac{CD-BE}{a}=\dfrac{(y-x)(CD-BE)}{a(CD+BE-2x)}"

"CD+BE-2x=y-x"

"CD+BE-2x=y+x"

Area of the trapezoid "BCDE"


"A(x, y)=\dfrac{x+y}{2}\cdot a \ \ ({units}^2)"